Plaese help!

Suppose you have an infinite non-conducting cylind
Answer #1

consider a long cylinder of radius R and uniform charge density \lambda let us find the electric field at a distance r from the axis cylinder

consider a cylindrical gaussian srface of length L and radius r

basing on electric flux

\int E.ds=Qenclosed/\epsilon

Qenclosed =\lambda r L

E[2\pi R*L]=\lambda rL/\epsilon R

E=\lambda r/2 \pi \epsilon R^2

[b] ans

basing on gauss law

\int E.ds=Qenclosed/\epsilon

Qenclosed =\lambda L

E[2\pi rL]=\lambda L/\epsilon

E=\lambda /2\pi \epsilon r

hence the therom is proved